# On subgroups with narrow Schreier graphs

### v2

@@ 1 Introduction
The number of relative ends of a subgroup $H\leq G$ was introduced by Houghton in the same paper. It can be defined as follows. Fix a finite generating set $S$ for $G$ and let $\operatorname{Cay}(G,S)$ be the Cayley graph of $G$ with respect to $S$. Given a subspace $K\subseteq\operatorname{Cay}(G,S)$, we say that a connected component of $\operatorname{Cay}(G,S)-K$ is deep if it contains points arbitrarily far from $K$. Given $R\geq 0$, let $\mathcal{N}_{R}(H)$ denote the $R$-neighbourhood of $H$. The number of relative ends of $H$ is the maximum over $R\geq 0$ of the number of $H$-orbits of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. In contrast, the number of filtered ends of $H$ is the maximum over $R\geq 0$ of the number of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. This invariant was introduced by Kropholler and Roller in [KR89]; a more geometric perspective can be found in [Geo08, Chapter 14].
@@
– is separable if for every $g\in G-H$ there is a finite index subgroup $K\leq G$ such that $H\leq K$ and $g\notin K$;
@@
The descending implications of the theorem all hold without any assumptions on $H\operatorname{\setminus}G$. The implications (1) $\Rightarrow$ (2) and (3) $\Rightarrow$ (4) are straightforward while (2) $\Rightarrow$ (3) is due to Yang [Yan11]. In contrast, none of the ascending implications hold without some extra condition on $H$ or on $G$ (see Examples 1.3, 1.4, 1.5 below). Moreover, none of the properties (1)-(4) follow directly from the fact that $H\operatorname{\setminus}G$ is narrow with at least two ends, or even from the stronger assumption that $H\operatorname{\setminus}G$ is quasi-isometric to a line. This is illustrated in Section 6 using a class of groups introduced by Houghton [Hou78]. Finally, the necessity of the condition that $H$ is finitely generated is illustrated by Example 1.6 below.
@@
Let $G$ be the Grigorchuk group – viewed as a group of automorphisms of the 2-regular rooted tree $T$ – and recall that $G$ is a finitely generated torsion group so in particular cannot contain any virtual fiber subgroups (see e.g. [Gri05]). Given a ray $\xi$ based at the root of $T$, let $G_{\xi}$ denote the stabiliser of $\xi$. The stabiliser of each vertex of $\xi$ has finite index in $G$ so $G_{\xi}$ is an intersection of finite index subgroups and is therefore separable. For many choices of $\xi$, the Schreier graph of $G_{\xi}$ is quasi-isometric to a line (see e.g. [Gri05, page 201]) so in particular it is narrow and two-ended yet $G_{\xi}$ does not satisfy the conclusion of Theorem 1.1.
The following corollary is an immediate consequence of Theorem 1.1 and Proposition 3.2.
@@
Proof.
Suppose that $H$ is a virtual fiber subgroup and let $H^{\prime}\leq H$ be such that the normaliser $N(H^{\prime})$ has finite index in $G$ and $N(H^{\prime})/H^{\prime}=\mathbb{Z}$. Then $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$ is a quasi-line and there is a covering map $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)\rightarrow H\operatorname{\setminus}\operatorname{Cay}(G,S)$. If $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ has $\geq n$ ends then there are $n$ geodesic rays $\overline{\delta_{1}},\dots,\overline{\delta_{n}}\subseteq H\operatorname{\setminus}\operatorname{Cay}(G,S)$ with infinite Hausdorff distance from each other. If $\delta_{1},\dots,\delta_{n}\subseteq H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$ are lifts of $\overline{\delta_{1}},\dots,\overline{\delta_{n}}$ respectively then for all $i\neq j$, $\delta_{i}$ and $\delta_{j}$ have infinite Hausdorff distance from each other. Since $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$ is a quasi-line, this implies that $n\in\{1,2\}$. Thus the corollary follows from Theorem 1.1. ∎
@@
Let $\operatorname{Sub}(G)$ denote the set of all subgroups of $G$ equipped with the topology induced from the Cantor set $\{0,1\}^{G}$ (viewed as the power set of $G$). In previous work with Damien Gaboriau, we showed that if $G$ is finitely generated and $H\leq G$ is such that $H\operatorname{\setminus}G$ is quasi-isometric to a line then there are countably many intermediate subgroups $H\leq K\leq G$ [AG23, Proposition 15]. In the case where $H$ is finitely generated this implies that $H$ is not in the perfect kernel of $\operatorname{Sub}(G)$. More precisely, we showed that such a subgroup $H$ vanishes by the third Cantor-Bendixson derivative of $\operatorname{Sub}(G)$ [AG23, Theorem 3.10] (i.e. $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$ in the language of that paper). Combining this with the results of this paper we obtain the following.
@@
Let $H\leq K\leq G$ be an intermediate subgroup. By Proposition 3.5 either $[G:K]<\infty$ or $[K:H]<\infty$. For each $k\in\mathbb{N}$, there are finitely many subgroups of $G$ of index $k$ and there are at most countably many subgroups of $G$ containing $H$ with index $k$ so this implies that there are countably many such subgroups $H\leq K\leq G$. Suppose that $H$ is finitely generated and that $H\operatorname{\setminus}G$ has at least two ends. Theorem 1.1 implies that either $H$ is a virtual fiber subgroup or $|H\operatorname{\setminus}G/H|<\infty$. In the former case $H\operatorname{\setminus}G$ is a quasi-line so $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$ by [AG23, Theorem 3.10] and in the latter $H$ is isolated in $\operatorname{Sub}(G)$ so $\operatorname{rk}_{CB}^{\times}(G;H)=1$. ∎
@@
The main content of Theorem 1.1 is the implication $(3)\Rightarrow(1)$ and this is proved using an action of $G$ on a finite dimensional CAT(0) cube complex. This cube complex is obtained using Sageev’s construction [Sag95] applied to the codimension one subgroup $H$ and the fact that it is finite dimensional follows from the fact that $H$ has bounded packing [HW09]. This will allow us to use techniques developed by Caprace and Sageev in [CS11], some of which we recall in the next section. We will conclude using the following proposition, which may be of independent interest.
@@ Acknowledgements.
I am very grateful to Mark Hagen and Indira Chatterji for many helpful comments and suggestions and for carefully reading several drafts of this paper. I thank Yves Cornulier and Paul-Henry Leemann for pointing out Example 1.6 and Adrien Le Boudec for asking me whether a previous version of Theorem 1.1 could be extended to include subgroups whose Schreier graphs have linear growth. This work was funded by a University of Bristol PhD scholarship.
@@ 2 Cube complexes and facing triples
We assume some familiarity with actions on CAT(0) cube complexes; introductory material can be found in [CN05, HW14, Nic04]. Recall that in [Sag95] Sageev gave a construction which, given a pair of groups $H\leq G$ such that $H$ has at least two relative ends, produces a CAT(0) cube complex on which $G$ acts with a single orbit of hyperplanes such that $H$ has finite index in the stabiliser of some hyperplane. As was pointed out by Haglund and Wise in [HW08, Corollary 3.1], it was observed in [Sag97] that if $H$ is finitely generated and has bounded packing then this construction yields a finite dimensional cube complex (see also [HW14, Corollary 3.31]).
@@
We will work with the piecewise $\ell^{1}$ metric on CAT(0) cube complexes (see [Mie14]) which we denote by $d$. Let $g$ be a cubical automorphism of a CAT(0) cube complex $C$ and recall that $g$ is an isometry with respect to $d$ and the CAT(0) metric (i.e. the piecewise $\ell^{2}$ metric). A combinatorial geodesic in $C$ is a geodesic path $\ell:[0,L]\rightarrow C^{(1)}$ where $L\in\mathbb{N}$ and $\ell(i)\in C^{(0)}$ for each $i\in\mathbb{Z}\cap[0,L]$. Haglund showed in [Hag23, Theorem 6.3] that $g$ either fixes a point in $C$ or admits a combinatorial geodesic axis in the cubical subdivision $C^{\prime}$ of $C$ (see [Hag23, Definition 2.3, Lemma 4.2]). In the later case, $g$ is called hyperbolic and admits a (not necessarily unique or combinatorial) geodesic axis in $C$ (i.e. a bi-infinite geodesic on which $\operatorname{\langle}g\operatorname{\rangle}$ acts by translation).
@@
Let us recall some notions and results from [CS11]. Let $C$ be a finite dimensional cube complex and $G$ be a group acting by cubical automorphisms on $C$. The action is said to be essential if one (equivalently every) $G$-orbit contains points arbitrarily far from any halfspace of $G$. If $\operatorname{\mathfrak{h}}$ is a hyperplane of $C$ then we denote by $\operatorname{\mathfrak{h}}^{+},\operatorname{\mathfrak{h}}^{-}$ the halfspaces bounded by $\operatorname{\mathfrak{h}}$. For any $g\in G$ we write $g\operatorname{\mathfrak{h}}^{+}$ (respectively $g\operatorname{\mathfrak{h}}^{-}$) to denote the image of $\operatorname{\mathfrak{h}}^{+}$ (respectively $\operatorname{\mathfrak{h}}^{-}$) under $g$ and $(g\operatorname{\mathfrak{h}})^{+}$ to denote an independent choice of halfspace bounded by $g\operatorname{\mathfrak{h}}$. An element $g\in G$ is said to skewer $\operatorname{\mathfrak{h}}$ if there is some $n\in\mathbb{N}$ such that either $g^{n}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ or $g^{n}\operatorname{\mathfrak{h}}^{-}\subsetneq\operatorname{\mathfrak{h}}^{-}$. Caprace and Sageev show in [CS11, Lemma 2.3] that a element $g\in G$ skewers a hyperplane $\operatorname{\mathfrak{h}}$ if and only if $g$ is hyperbolic and one (equivalently any) axis of $g$ intersects $\operatorname{\mathfrak{h}}$ at a single point. This statement refers to geodesic axes of $g$ with respect to the CAT(0) metric but the same proof shows that the statement holds for $\ell^{1}$ geodesic axes of $g$, including combinatorial axes when they exist. The following characterisation of essentiality will be useful.
@@ 3 Narrow spaces
Now suppose that (b) holds. Let $e\in\mathbb{N}$ be the number of ends of $Y$. Fix $\mu\geq 1$ and let $D\geq\mu$ be such that any ball of radius $D$ separates $Y$ into at least two unbounded connected components. Let $R\geq 1$ be such that any ball of radius $D$ in $Y$ contains at most $R$ vertices and let $L\coloneqq eR$. Let $\xi_{1},\dots,\xi_{e}$ denote the ends of $Y$. Suppose that $Y_{1},\dots,Y_{L+1}\subseteq Y$ are unbounded $\mu$-coarsely connected subspaces, ordered such that the following holds. There exists $21=L_{0}\leq L_{1}\leq\dots\leq L_{e-1}\leq L_{e}=L+$ 2 such that, for each $i\in\{1,\dots,e\}$, if $j\in[L_{i-1},L_{i})$, then there is a subset of $Y_{j}$ which converges towards $\xi_{i}$. It follows that, for each $i\in\{1,\dots,e\}$, there is a ball $B_{i}\subseteq Y$ of radius $D$ which intersects $Y_{j}$ for all $j\in[L_{i-1},L_{i})$ Since the number of the vertices in $\cup_{i=1}^{e}B_{i}$ is at most $L$, there exists $j\neq k$ such that $Y_{j}\cap Y_{k}\neq\emptyset$. ∎
@@
Let $S$ be a finite generating set for $G$, let $X_{H}\coloneqq H\operatorname{\setminus}\operatorname{Cay}(G,S)$, $X_{K}\coloneqq K\operatorname{\setminus}\operatorname{Cay}(G,S)$ and let $\rho_{H}:\operatorname{Cay}(G,S)\rightarrow X_{H}$, $\rho_{K}:\operatorname{Cay}(G,S)\rightarrow X_{K}$ denote the quotient maps. Note that $\rho_{H}$ and $\rho_{K}$ are covers and that $\rho_{K}$ factors through a covering map $\pi:X_{H}\rightarrow X_{K}$ whose degree is $[K:H]$. Suppose that $[G:K]=\infty$, fix $\overline{x}\in X_{K}$ and let $\overline{\delta}\subseteq X_{K}$ be an infinite geodesic ray based at $\overline{x}$. For each point $x\in\pi^{-1}(\overline{x})$ there is a lift of $\overline{\delta}$ based at $x$. These lifts are pairwise disjoint and connected so $|\pi^{-1}(\overline{x})|\leq L(1)$, where $L(1)$ is as in Definition 3.1. Thus $[K:H]\leq L(1)<\infty$. ∎
@@
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $X_{H}$ is narrow. If $H$ has bounded packing then $H$ is a virtual fiber subgroup.

### v3

@@ 1 Introduction
The number of relative ends of a subgroup $H\leq G$ was introduced by Houghton in the same paper. It can be defined as follows. Fix a finite generating set $S$ for $G$ and let $\operatorname{Cay}(G,S)$ be the Cayley graph of $G$ with respect to $S$. Given a subspace $K\subseteq\operatorname{Cay}(G,S)$, we say that a connected component of $\operatorname{Cay}(G,S)-K$ is deep if it contains points arbitrarily far from $K$. Given $R\geq 0$, let $\mathcal{N}_{R}(H)$ denote the $R$-neighbourhood of $H$. The number of relative ends of $H$ is the supremum over $R\geq 0$ of the number of $H$-orbits of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. In contrast, the number of filtered ends of $H$ is the supremum over $R\geq 0$ of the number of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. This invariant was introduced by Kropholler and Roller in [KR89]; a more geometric perspective can be found in [Geo08, Chapter 14].
@@
– is separable if for every $g\in G-H$ there is a finite index subgroup $K\leq G$ such that $H\leq K$ and $g\notin K$, or equivalently if $H$ is closed with respect to the profinite topology on $G$;
@@
The descending implications of the theorem all hold if we only assume that $G$ is finitely generated and $[G:H]=\infty$. The implication (2) $\Rightarrow$ (3) is due to Yang [Yan11] and the other two are straight-forward. To deduce (1) $\Rightarrow$ (2), suppose that $H$ is a virtual fiber subgroup and let $H^{\prime}\leq H$ and $G^{\prime}\leq G$ be finite index subgroups such that there is a surjective homomorphism $\pi:G^{\prime}\rightarrow\mathbb{Z}$ with kernel $H^{\prime}$. Then $H^{\prime}=\cap_{n\in\mathbb{N}}\pi^{-1}(n\mathbb{Z})$ and each $\pi^{-1}(n\mathbb{Z})$ has finite index in $G$. Therefore $H^{\prime}$ is closed with respect to the profinite topology on $G$ and, since $H$ is a finite union of cosets of $H^{\prime}$, so is $H$. To see that (3) $\Rightarrow$ (4), note that $d_{S}(H,hgH)=d_{S}(H,gH)$ for all $g\in G$ and $h\in H$. Thus if $|H\operatorname{\setminus}G/H|<\infty$ then $d_{S}(gH,g^{\prime}H)=d_{S}(H,g^{-1}g^{\prime}H)$ is uniformly bounded and $H$ cannot have bounded packing unless $[G:H]<\infty$.
In contrast, none of the ascending implications hold without some extra condition on $H$ or on $G$ (see Examples 1.3, 1.4, 1.5 below). Moreover, none of the properties (1)-(4) follow directly from the fact that $H\operatorname{\setminus}G$ is narrow with at least two ends, or even from the stronger assumption that $H\operatorname{\setminus}G$ is quasi-isometric to a line. This is illustrated in Section 6 using a class of groups introduced by Houghton [Hou78]. Finally, the necessity of the condition that $H$ is finitely generated is illustrated by Example 1.6 below.
@@
Let $G$ be the Grigorchuk group – viewed as a group of automorphisms of the 2-regular rooted tree $T$ – and recall that $G$ is a finitely generated torsion group, so in particular cannot contain any virtual fiber subgroups (see e.g. [Gri05]). Given a ray $\xi$ based at the root of $T$, let $G_{\xi}$ denote the stabiliser of $\xi$. The stabiliser of each vertex of $\xi$ has finite index in $G$ so $G_{\xi}$ is an intersection of finite index subgroups and is therefore separable. For many choices of $\xi$, the Schreier graph of $G_{\xi}$ is quasi-isometric to a line (see e.g. [Gri05, page 201]) so in particular it is narrow and two-ended yet $G_{\xi}$ does not satisfy the conclusion of Theorem 1.1.
There are many examples of group pairs that satisfy the conditions of Theorem 1.1. For instance if $H\unlhd G$ is a finitely generated pair such that $G/H=\mathbb{Z}$ then $G/H$ is a quasi-line, so it is narrow by Proposition 3.2 and has two ends. More generally, if $H$ is a virtual fiber subgroup of a finitely generated group $G$ then $H\operatorname{\setminus}G$ is quasi-isometric to either $\mathbb{Z}$ or $\mathbb{N}$ (see Corollary 3.3) so Theorem 1.1 is close to optimal. This raises the following natural question.
Question 1.7.
Does there exist a finitely generated pair of groups $H\leq G$ such that $H\operatorname{\setminus}G$ is narrow and one-ended (e.g. a quasi-ray) and $|H\operatorname{\setminus}G/H|=\infty$ but $H$ is not a virtual fiber subgroup?
The following corollary follows from Theorem 1.1 and Proposition 3.2.
@@
Let $\operatorname{Sub}(G)$ denote the set of all subgroups of $G$ equipped with the topology induced from the Cantor set $\{0,1\}^{G}$ (viewed as the power set of $G$). In previous work with Damien Gaboriau, we showed that if $G$ is finitely generated and $H\leq G$ is such that $H\operatorname{\setminus}G$ is quasi-isometric to a line then there are countably many intermediate subgroups $H\leq K\leq G$ [AG23, Proposition 3.15]. In the case where $H$ is finitely generated this implies that $H$ is not in the perfect kernel of $\operatorname{Sub}(G)$. More precisely, we showed that such a subgroup $H$ vanishes by the third Cantor-Bendixson derivative of $\operatorname{Sub}(G)$ [AG23, Theorem 3.10] (i.e. $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$ in the language of that paper). Combining this with the results of this paper we obtain the following.
@@
Let $H\leq K\leq G$ be an intermediate subgroup. By Proposition 3.6 either $[G:K]<\infty$ or $[K:H]<\infty$. For each $k\in\mathbb{N}$, there are finitely many subgroups of $G$ of index $k$ and there are at most countably many subgroups of $G$ containing $H$ with index $k$ so this implies that there are countably many such subgroups $H\leq K\leq G$. Suppose that $H$ is finitely generated and that $H\operatorname{\setminus}G$ has at least two ends. Theorem 1.1 implies that either $H$ is a virtual fiber subgroup or $|H\operatorname{\setminus}G/H|<\infty$. In the former case Corollary 3.3 implies that $H\operatorname{\setminus}G$ is a quasi-line so $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$ by [AG23, Theorem 3.10] and in the latter $H$ is isolated in $\operatorname{Sub}(G)$ so $\operatorname{rk}_{CB}^{\times}(G;H)=1$. ∎
@@
The main content of Theorem 1.1 is the implication $(3)\Rightarrow(1)$ and this is proved using an action of $G$ on a finite dimensional CAT(0) cube complex $C$. This cube complex is obtained using Sageev’s construction [Sag95] applied to the codimension one subgroup $H$ and the fact that it is finite dimensional follows from the fact that $H$ has bounded packing [HW09]. This will allow us to use tools developed by Caprace and Sageev in [CS11] to show that $C$ has no facing triples of hyperplanes. We will conclude using the following proposition, which may be of independent interest.
@@ Acknowledgements.
I am very grateful to Mark Hagen and Indira Chatterji for many helpful comments and suggestions and for carefully reading several drafts of this paper. I thank Yves Cornulier and Paul-Henry Leemann for pointing out Example 1.6 and Adrien Le Boudec for asking me whether a previous version of Theorem 1.1 could be extended to include subgroups whose Schreier graphs have linear growth. I am grateful to the anonymous referee for a number of helpful comments and corrections. This work was funded by a University of Bristol PhD scholarship.
@@ 2 Cube complexes and facing triples
We assume some familiarity with actions on CAT(0) cube complexes; introductory material can be found in [CN05, HW14, Nic04]. Recall that in [Sag95] Sageev gave a construction which, given a pair of groups $H\leq G$ such that $H$ has at least two relative ends, produces an unbounded CAT(0) cube complex on which $G$ acts with a single orbit of hyperplanes such that $H$ has finite index in the stabiliser of some hyperplane $\mathfrak{h}$. As was pointed out by Haglund and Wise in [HW08, Corollary 3.1], it was observed in [Sag97] that if $H$ is finitely generated and has bounded packing then Sageev’s construction yields a finite dimensional cube complex (see also [HW14, Corollary 3.31]).
@@
We will work with the piecewise $\ell^{1}$ metric on CAT(0) cube complexes (see [Mie14]) which we denote by $d$. Let $g$ be a cubical automorphism of a CAT(0) cube complex $C$ and recall that $g$ is an isometry with respect to both $d$ and the CAT(0) metric (i.e. the piecewise $\ell^{2}$ metric). A combinatorial geodesic in $C$ is a geodesic path $\ell:[0,L]\rightarrow C^{(1)}$ where $L\in\mathbb{N}$ and $\ell(i)\in C^{(0)}$ for each $i\in\mathbb{Z}\cap[0,L]$. Haglund showed in [Hag23, Theorem 6.3] that $g$ either fixes a point in $C$ or admits a combinatorial geodesic axis in the cubical subdivision $C^{\prime}$ of $C$ (see [Hag23, Definition 2.3, Lemma 4.2]). In the later case, $g$ is called hyperbolic and admits a (not necessarily unique or combinatorial) geodesic axis in $C$ (i.e. a bi-infinite $\ell^{1}$-geodesic on which $\operatorname{\langle}g\operatorname{\rangle}$ acts by translation). Such an isometry $g$ is also hyperbolic in the usual CAT(0) sense and as such also admits an $\ell^{2}$ geodesic axis.
@@
Let us recall some notions and results from [CS11]. Let $C$ be a finite dimensional cube complex and $G$ be a group acting by cubical automorphisms on $C$. The action is said to be essential if one (equivalently every) $G$-orbit contains points arbitrarily far from any halfspace of $G$. If $\operatorname{\mathfrak{h}}$ is a hyperplane of $C$ then we denote by $\operatorname{\mathfrak{h}}^{+},\operatorname{\mathfrak{h}}^{-}$ the halfspaces bounded by $\operatorname{\mathfrak{h}}$. For any $g\in G$ we write $g\operatorname{\mathfrak{h}}^{+}$ (respectively $g\operatorname{\mathfrak{h}}^{-}$) to denote the image of $\operatorname{\mathfrak{h}}^{+}$ (respectively $\operatorname{\mathfrak{h}}^{-}$) under $g$ and $(g\operatorname{\mathfrak{h}})^{+}$ to denote an independent choice of halfspace bounded by $g\operatorname{\mathfrak{h}}$. An element $g\in G$ is said to skewer $\operatorname{\mathfrak{h}}$ if there is some $n\in\mathbb{N}$ such that either $g^{n}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ or $g^{n}\operatorname{\mathfrak{h}}^{-}\subsetneq\operatorname{\mathfrak{h}}^{-}$. Caprace and Sageev show in [CS11, Lemma 2.3] that an element $g\in G$ skewers a hyperplane $\operatorname{\mathfrak{h}}$ if and only if $g$ is hyperbolic and one (equivalently any) axis of $g$ intersects $\operatorname{\mathfrak{h}}$ at a single point. This statement refers to geodesic axes of $g$ with respect to the CAT(0) metric but the same proof shows that the statement holds for combinatorial geodesic axes of $g$. The following characterisation of essentiality will be useful.
@@ 3 Narrow spaces
Now suppose that (b) holds and let $T$ be a finitely ended tree and $f:T\rightarrow Y$ be a $\lambda$-quasi-isometry for some $\lambda\geq 1$. Let $D=3\lambda^{3}+\lambda^{2}+3\lambda$ and let $\overline{f}:T\cup\partial T\rightarrow Y\cup\partial Y$ denote the extension of $f$ to the visual compactification of $T$. We henceforth identify the ends of $T$ with the points in its visual boundary in the natural way. Let $e\in\mathbb{N}$ be the number of ends of $T$ and fix $\mu\geq 1$. Let $R\geq 1$ be such that any ball of radius $D+\mu$ in $Y$ contains at most $R$ vertices and let $L\coloneqq eR$. Let $\xi_{1},\dots,\xi_{e}$ denote the ends of $T$ and note that $\overline{f}(\xi_{1}),\dots\overline{f}(\xi_{e})$ are the ends of $Y$. Suppose that $Y_{1},\dots,Y_{L+1}\subseteq Y$ are unbounded $\mu$-coarsely connected subspaces, ordered such that the following holds. There exists $21=L_{0}\leq L_{1}\leq\dots\leq L_{e-1}\leq L_{e}=L+$ 2 such that, for each $i\in\{1,\dots,e\}$, if $j\in[L_{i-1},L_{i})$, then there is a subset of $Y_{j}$ which converges towards $\overline{f}(\xi_{i})$. For each $i\in\{1,\dots,e\}$ let $\gamma_{i}\subseteq T$ be a geodesic ray which is a representative of $\xi_{i}$. Up to removing a finite length segment from $\gamma_{i}$, we can assume that each vertex of $\gamma_{i}$ separates $T$ into at least two unbounded connected components. It follows that for each vertex $v\in\gamma_{i}$, the ball of radius $D$ around $f(v)$ separates $Y$ into at least two unbounded connected components. It follows that, if $j\in[L_{i-1},L_{i})$, then, for all but finitely many $v\in\gamma_{i}$, the subspace $Y_{j}$ contains vertices lying in at least two distinct connected components of $Y-B_{Y}(f(v),D)$. Therefore there exists a vertex $v_{i}\in Y$ such that, for all $j\in[L_{i-1},L_{i})$, the subspace $Y_{j}$ contains vertices lying in at least two distinct connected components of $Y-B_{Y}(v_{i},D)$. If $B_{Y}(v_{i},D)\cap Y_{j}=\emptyset$ then there are vertices $x,y\in Y_{j}$ lying in different connected components of $Y-B_{Y}(v_{i},D)$ such that $d(x,y)\leq\mu$. Let $z\in B_{Y}(v_{i},D)$ be a vertex which lies on a geodesic from $x$ to $y$. Then $d(v_{i},x)\leq d(v_{i},z)+d(z,x)<D+\mu$. Therefore $Y_{j}\cap B_{Y}(v_{i},D+\mu)\neq\emptyset$ for all $j\in[L_{i-1},L_{i})$. Since the number of the vertices in $\cup_{i=1}^{e}B_{Y}(v_{i},D+\mu)$ is at most $L$, there exists $j\neq k$ such that $Y_{j}\cap Y_{k}\neq\emptyset$. ∎
Corollary 3.3.
Let $G$ be a finitely generated group and let $H\leq G$ be a virtual fiber subgroup. Let $S$ be any finite generating set of $G$. Then $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ is quasi-isometric to either $\mathbb{Z}$ or $\mathbb{N}$. In particular $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ is narrow.
Proof.
Let $H^{\prime}\leq H$ and $G^{\prime}\leq G$ be finite index subgroups such that $H^{\prime}\unlhd G^{\prime}$ and $G^{\prime}/H^{\prime}=\mathbb{Z}$. Let $X_{H}\coloneqq H\operatorname{\setminus}\operatorname{Cay}(G,S)$ and $X_{H^{\prime}}\coloneqq H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$. Then, for any finite generating set $S^{\prime}$ of $G^{\prime}$, the graph $X_{H^{\prime}}$ is quasi-isometric to $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G^{\prime},S^{\prime})$, which is a quasi-line. Moreover the quotient map $\rho:X_{H^{\prime}}\rightarrow X_{H}$ defined by $\rho(H^{\prime}g)=Hg$ is a finite degree covering map. Suppose that there exists $K\in\mathbb{N}$ such that $\operatorname{Diam}\rho^{-1}(x)\leq K$ for all $x\in X_{H}$. Let $x,y\in X_{H^{\prime}}$ and let $\gamma\subseteq X_{H}$ be a geodesic from $\rho(x)$ to $\rho(y)$. Let $y^{\prime}\in X_{H^{\prime}}$ be the endpoint of the lift of $\gamma$ based at $x$. Then $d_{X_{H^{\prime}}}(x,y)\leq d_{X_{H^{\prime}}}(x,y^{\prime})+d_{X_{H^{\prime}}}(y^{\prime},y)\leq d_{X_{H}}(\rho(x),\rho(y))+K$. Since $\rho$ is a surjective 1-Lipschitz map, this implies that $\rho$ is a quasi-isometry. If $\operatorname{Diam}\rho^{-1}(x)$ is not uniformly bounded then [AG23, Lemma 3.14] implies that there is a geodesic ray $\gamma\subseteq X_{H}$ and two lifts $\gamma_{1},\gamma_{2}\subseteq X_{H^{\prime}}$ of $\gamma$ such that the Hausdorff distance between $\gamma_{1}$ and $\gamma_{2}$ is infinite. Then $X_{H^{\prime}}$ is quasi-isometric to the union $\gamma_{1}\cup\gamma_{2}$ (equipped with the induced metric from $X_{H^{\prime}}$) and it follows that $X_{H}$ is quasi-isometric to $\gamma$. Proposition 3.2 then implies that $X_{H}$ is narrow. ∎
@@
Let $S$ be a finite generating set for $G$, let $X_{H}\coloneqq H\operatorname{\setminus}\operatorname{Cay}(G,S)$, $X_{K}\coloneqq K\operatorname{\setminus}\operatorname{Cay}(G,S)$ and let $\rho_{H}:\operatorname{Cay}(G,S)\rightarrow X_{H}$, $\rho_{K}:\operatorname{Cay}(G,S)\rightarrow X_{K}$ denote the quotient maps. Note that $\rho_{H}$ and $\rho_{K}$ are covers and that $\rho_{K}$ factors through a covering map $\pi:X_{H}\rightarrow X_{K}$ whose degree is $[K:H]$. Suppose that $[G:K]=\infty$, fix $\overline{x}\in X_{K}$ and let $\overline{\delta}\subseteq X_{K}$ be an infinite geodesic ray based at $\overline{x}$. For each point $x\in\pi^{-1}(\overline{x})$ there is a lift of $\overline{\delta}$ based at $x$. These lifts are pairwise disjoint and connected so $|\pi^{-1}(\overline{x})|\leq L(1)$, where $L:\mathbb{N}\rightarrow\mathbb{N}$ is a map which expresses the narrowness of $X_{H}$ (see Definition 3.1). Thus $[K:H]\leq L(1)<\infty$. ∎
@@
Let $G$ be a group with finite generating set $S$ and $H\leq G$ be a finitely generated subgroup such that $X_{H}$ is narrow and has at least two ends. If $H$ has bounded packing then $H$ is a virtual fiber subgroup.